OpenStudy (anonymous):

Explain the connections and parallels in finding missing side lengths and angle measures in a right triangle using the trigonometric functions.

OpenStudy (anonymous):

u can use sin cos tan formula

OpenStudy (shinalcantara):



OpenStudy (shinalcantara):

REMEMBER: \[\cos \theta = \frac{ adjacent side }{ hypotenuse }\] \[\sin \theta = \frac{ opposite side }{ hypotenuse }\] \[\tan \theta = \frac{ opposite side }{ adjacent side }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (shinalcantara):

yw :)